OpenStudy (anonymous):

If the cos 90° = 0, then the sin 0° = _____.

OpenStudy (johnweldon1993):

Cos 90 is indeed 0 Sin 0 is Also 0

OpenStudy (anonymous):

0?

zepdrix (zepdrix):

Sine and Cosine are cofunctions:\[\Large\bf\sf \cos(\text{angle})\quad=\quad \sin(90^o-\text{angle})\] \[\Large\bf\sf \cos(90^o)\quad=\quad \sin(90^o-90^o)\]